OpenStudy (anonymous):

xb-ab+ax-x^2 Factor Completely I get b(x-a)+x(a-x) = (b+x) then what? How am I messing up?

OpenStudy (anonymous):

The factorisation can be as follows: -x^2+xb+ax-ab = x(-x + b) -a( -x +b) = (x-a) (b-x)

OpenStudy (anonymous):

oh yeah thank youuuuu